OpenStudy (anonymous):

A dog with mass 14 kg has a kinetic energy of 135.5 J. What is the speed of the dog? A. 3.2 m/s B. 2.3 m/s C. 4.4 m/s D. 5.1 m/s

OpenStudy (anonymous):

Well \[E = \frac{ 1 }{ 2 }mv^2\] Rearrange to find v, \[v = \sqrt{\frac{2E }{ m }}\] \[v = \sqrt{\frac{ 2(135.5) }{ 14 }}\] = 4.4m/s